OpenStudy (anonymous):

(3*x^2-6*x-24)/(5*x^2-26*x+5)

OpenStudy (anonymous):

3x^2-6x-24=3x^2-12x+6x-24=3x(x-4)+6(x-4)=(3x+6)(x-4) 5x^2-26x+5=5x^2-25x-x+5=5x(x-5)-1(x-5)=(5x-1)(x-5) ans=(3x+6)(x-4)/(5x-1)(x-5)